# Mixstyle based Domain Generalization for Sound Event Detection with Heterogeneous Training Data

### v1

@@ 3.3 Metrics
The DCASE 2024 Task 4 considers two metrics for evaluation. PSDS [28, 29] is computed based on event onset and offset times, which is only available for DESED data and thereby only on that fraction of the evaluation set. For MAESTRO, segment-based labels (one second) are provided, and the segment-based mean (macro-averaged) partial area under the ROC curve (mpAUC) is used as the primary metric, with a maximum FP-rate of 0.1. mpAUC is computed with respect to hard labels (threshold = 0.5) for the 11 classes listed. It is noted that the DESED and MAESTRO clips are anonymized and shuffled in the evaluation set to prevent manual domain identification.
@@ 4 Result and Analysis
We are first interested in evaluating the impact of the proposed DG method used on DESED dataset in terms of PSDS metric that measures how well the system detects the sound events. From Table 1, we can observe that the baseline system has a PSDS of 0.549 on the public evaluation set. Introducing frequency-dynamic convolution (FDY) improves the performance by enhancing generalization. Then, Freq-MixStyle further increases the performance by improving data diversity and robustness. By combining FDY and Freq-MixStyle, we obtain a better result, especially with cSEBBs post-processing. Further, on applying AdaResNorm to Freq-MixStyle helps with normalization to improve the PSDS. The best PSDS of 0.604 on the public evaluation set is obtained when all the above methods are used, benefits from each method.
Then we analyze the results in terms of mpAUC, which measures performance on the MAESTRO dataset as shown in Table 1. Similar to the trend we observed in case of the DESED dataset, each method shows their contribution to enhance the baseline performance on MAESTRO dataset. On introducing FDY, Freq-MixStyle and AdaResNorm together, we obtain the best mpAUC of 0.739 on MAESTRO Real validation set. This confirms our proposed DG method works effectively on both DESED and MAESTRO datasets.
Finally, we compute the joint score that combines PubEval-PSDS and mpAUC scores for each system as DCASE 2024 Challenge Task 4 considers the joint score for benchmarking the systems. From Table 1, we observe that the joint score significantly improves from 1.270 to 1.343 on introducing FDY, Freq-MixStyle, AdaResNorm and cSEBBs to the baseline system, showcasing their contributions. While FDY provides a larger improvement compared to Freq-MixStyle (especially on DESED subset), it comes at the cost of increase in model parameters around twice that in the baseline. On the contrary, Freq-MixStyle is useful for developing low-complexity systems as it does not affect on the model parameters.
@@ 5 Conclusion
This work demonstrates the effectiveness of DG-based techniques for SED systems trained with heterogeneous data. By integrating Freq-MixStyle, adaptive residual normalization, and cSEBBs within a mean-teacher framework, our approach successfully adapts mel-spectrograms and generalizes features across multiple domains. Studies on the DCASE 2024 Challenge Task 4 dataset shows that our DG-based method significantly improves both the polyphonic SED score on the DESED dataset and the macro-average pAUC on the MAESTRO Real dataset, outperforming the challenge baseline. These results highlight the potential of DG-based methods in advancing SED adaptability to real-world scenarios.

### v3

@@ 3.3 Metrics
The DCASE 2024 Task 4 considers two metrics for evaluation. PSDS [34, 35] is computed based on event onset and offset times, which are “PSDS1” in the previous DCASE challenge and only available for DESED data, and thereby only on that fraction of the evaluation set. For MAESTRO, segment-based labels (one second) are provided, and the segment-based mean (macro-averaged) partial area under the ROC curve (mpAUC) is used as the primary metric, with a maximum FP-rate of 0.1. mpAUC is computed with respect to hard labels (threshold = 0.5) for the 11 classes listed. It is noted that the DESED and MAESTRO clips are anonymized and shuffled in the evaluation set to prevent manual domain identification.
@@ 4 Result and Analysis
We are first interested in evaluating the impact of the proposed DG-SED method used on the DESED dataset in terms of the PSDS metric that measures how well the system detects sound events. From Table 1, we can observe that the baseline CRNN system has a PSDS of 0.549 on the public evaluation set. And the FDY-CRNN [3] improves the performance with more parameters. DG-SED (without AdaResNorm) further increases the performance by improving data diversity for both CRNN and FDY-CRNN backbones. We obtain a better result with cSEBBs post-processing. Further, applying the whole DG-SED method helps with normalization to improve the PSDS. The best PSDS of 0.604 on the public evaluation set is obtained when all the modules are used, and benefits from each module.
Then we analyze the results in terms of mpAUC, which measures performance on the MAESTRO dataset as shown in Table 1. Similar to the trend we observed in case of the DESED dataset, each method shows its contribution to enhancing the baseline performance on the MAESTRO dataset. When introducing Freq-MixStyle, and AdaResNorm together as DG-SED to FDY-CRNN backbone, we obtain the best mpAUC of 0.739 on the MAESTRO Real validation set. This confirms our proposed DG-SED method works effectively on both DESED and MAESTRO datasets.
Finally, we compute the joint score that combines PubEval-PSDS and mpAUC scores for each system as DCASE 2024 Challenge Task 4 considers the joint score for benchmarking the systems. From Table 1, we observe that the joint score significantly improves from 1.270 to 1.314 upon introducing DG-SED with cSEBBs to the baseline CRNN system, showcasing their contributions. Also with the FDY-CRNN backbone, we obtain the best joint score, it comes at the cost of the increase in model parameters around twice that in the baseline. On the contrary, DG-SED is useful for developing low-complexity systems as it does not affect the model parameters.
This Table 2 shows that our DG-SED system, which incorporates domain generalization methods (e.g., MixStyle and Adaptive Residual Normalization) into the FDY-CRNN with cSEBBs, achieves a joint score of 1.343 on the DCASE 2024 Task 4 public evaluation set and . This score surpasses other top submissions under similar resource constraints. We avoid direct comparisons with models with very different parameter sizes, to ensure a fair evaluation. Our model requires only 3.4 million parameters, demonstrating that it enhances accuracy while maintaining a compact structure.
@@ 5 Conclusion
This work demonstrates the effectiveness of DG-based techniques for SED systems trained with heterogeneous data. By integrating Freq-MixStyle, adaptive residual normalization, and cSEBBs within a mean-teacher framework, our DG-SED approach successfully adapts mel-spectrograms and generalizes features across multiple domains. Studies on the DCASE 2024 Challenge Task 4 dataset show that our DG-SED method significantly improves both the polyphonic SED score on the DESED dataset and the macro-average pAUC on the MAESTRO Real dataset, outperforming the challenge baseline. These results highlight the potential of DG-SED methods in advancing SED adaptability to real-world scenarios.